geerky42 (geerky42):

What's difference between \(\dfrac{\text{d}}{\text{d}x}\) and \( \dfrac{∂}{∂x}\ \)?

OpenStudy (anonymous):

one is a partial derivative

OpenStudy (anonymous):

the first one is differentiation w.r.t x the second one is partial differentiation w.r.t. x

geerky42 (geerky42):

What's the difference? They seem like they work the same way. Exactly what's the difference?

OpenStudy (anonymous):

lets say f(x,y) = x^2* y

OpenStudy (anonymous):

partial diff w.r.t x, u get 2xy

OpenStudy (anonymous):

u take y as constant and diff w.r.t x

OpenStudy (anonymous):

if your a partial differentiating with respect to one variable in a multivariable problem, you treat the other variables as constants when differentiating.

OpenStudy (anonymous):

\[\frac{d}{dx}\] acts on a function of a single variable. Namely, x. \[\frac{\partial }{\partial x}\] acts on a function of more than one variable. For example: \[f(x,y,z)\]

geerky42 (geerky42):

What is "w.r.t?"

OpenStudy (anonymous):

if you diff w.r.t x, you will need to do implicit differentiation

OpenStudy (anonymous):

with respect to

OpenStudy (anonymous):

though, you cant do implicit differentiatin for a function that is f(x,y) partial implicit differentiation is fine

geerky42 (geerky42):

I see so differentiation treats other variable as variable and particle differentiation treats other variable as constant?

geerky42 (geerky42):

And in what situation will partial differentiation be useful?

OpenStudy (anonymous):

yes. but you cant use both on the same function. partial differentiation works for functions where the variables are not mutually exclusive

OpenStudy (anonymous):

as in f(x,y)

OpenStudy (anonymous):

d/dx is for functions where the variables are mutually exclusive. as in f(x) or f(y)

OpenStudy (anonymous):

you can d/dx an f(y) function but u cant d/dx an f(x,y) function

OpenStudy (anonymous):